Calculate the percent by mass of a NaCl solution with molar mass of 58.44g/mole placed in 15.5g of H2O with molar mass of 18.0g/mole.


1
Expert's answer
2021-03-22T06:17:43-0400

We will use 1 g as the value of NaCl mass.

Total mass of the solution = 1 + 15.5 = 16.5 g

The percent by mass of a NaCl "= \frac{1 \; g \times 100 \;\%}{16.5 \;g} = 6.06 \;\%"